# Connecting Discrete Morse Functions via Birth-Death Transitions

### v1

@@ 1. Introduction
This paper continues the study and clarifies some ambiguities in [17] by following the idea of birth-death transitions to connect discrete Morse functions on $\mathcal{M}(K)$. Precisely, we show in Theorem 4.6 that any two discrete Morse functions $f_{1}$ and $f_{2}$ can be transformed into one another with a finite sequence of birth-death transitions. Moreover, we apply this theorem to give an alternative, self-contained proof of some Forman’s fundamental results in Theorem 5.1 and Theorem 5.2.
@@ 3. Birth-death transitions between discrete Morse complexes
Suppose that $h:C_{q}^{f_{1}}\to C_{q}^{f_{2}}$ and $g:C_{q}^{f_{2}}\to C_{q}^{f_{1}}$ be a birth transition and a death transition, respectively.
Then, $h$ and $g$ are chain-homotopy inverses to each other, that is, $g\circ h$ and $h\circ g$ are chain homotopic to the identities.
@@
Therefore $g\circ h=\mathrm{id}$ and $h\circ g\simeq\mathrm{id}$, so $h$ and $g$ are chain-homotopy inverses. ∎
@@ 5. An alternative proof of Forman’s Theorems
One of the fundamental results in discrete Morse theory, established by Forman [7], is the equivalence between discrete Morse homology and simplicial homology. In our framework, where birth-death transitions act as chain maps between discrete Morse complexes, this correspondence can be recovered naturally.
@@
By Theorem 4.6, the gradient vector field $V$ corresponding to $f$ is connected to $V_{0}$ through a sequence of birth-death transitions, each of which defines a chain homotopy equivalence. Thus, composing the induced maps yields a homology isomorphism:
@@
Besides Theorem 5.1, other Forman’s theorems can also be recovered in the sense of “collapsing critical simplices”, similar to the proof above. We provide one more example, the Morse equality, as follows.

### v2

@@ 1. Introduction
This paper continues the study and clarifies some ambiguities in [17] by following the idea of birth-death transitions to connect discrete Morse functions on $\mathcal{M}(K)$. Precisely, we show in Theorem 4.6 that any two discrete Morse functions $f_{1}$ and $f_{2}$ can be transformed into one another with a finite sequence of birth-death transitions. Moreover, we apply this theorem to give an alternative, self-contained proof of some Forman’s fundamental results over $\mathbb{Q}$ in Theorem 5.1 and Theorem 5.2.
@@ 3. Birth-death transitions between discrete Morse complexes
Suppose $h:C_{\ast}^{f_{1}}\to C_{\ast}^{f_{2}}$ and $g:C_{\ast}^{f_{2}}\to C_{\ast}^{f_{1}}$ are a birth transition and a death transition, respectively.
Then, over $\mathbb{Q}$, $h$ and $g$ are chain-homotopy inverses to each other. That is, the induced chain maps $h\circ g$ and $g\circ h$ are chain homotopic to the identities on rational coefficients.
@@
Here $\tfrac{1}{k}\in\mathbb{Q}$, so this definition is valid over $\mathbb{Q}.$
@@
Therefore $g\circ h=\mathrm{id}$ and $h\circ g\simeq\mathrm{id}$, as chain maps over $\mathbb{Q}$, so $h$ and $g$ are chain-homotopy inverses in this setting.
See Remark 5.3 for more comments on coefficients.
@@ 5. An alternative proof of Forman’s Theorems
One of the fundamental results in discrete Morse theory, established by Forman [7], is the equivalence between discrete Morse homology and simplicial homology. In our framework, where birth-death transitions act as chain maps between discrete Morse complexes, this correspondence can be recovered over coefficient $\mathbb{Q}$ naturally.
@@
By Theorem 4.6, the gradient vector field $V$ corresponding to $f$ is connected to $V_{0}$ through a sequence of birth-death transitions, each of which defines a chain homotopy equivalence over $\mathbb{Q}$. Thus, composing the induced maps yields a homology isomorphism:
@@
Besides Theorem 5.1, other Forman’s theorems can also be recovered in the sense of “collapsing critical simplices” while keeping $K$ unchanged, similar to the proof above. We provide one more example, the Morse equality, as follows.
@@
Remark 5.3.
We make the following comments regarding coefficients:
Forman’s Theorems 5.1 and 5.2 are known to hold integrally, i.e. over $\mathbb{Z}$. Theorem 3.6 establishes that a birth–death transition induces a chain homotopy equivalence over $\mathbb{Q}$. The proof works integrally whenever the connectedness coefficient satisfies $|k|=1$, so in this case the statement holds over $\mathbb{Z}$, as well.
In smooth Cerf theory, bifurcations with $|k|\geq 2$ can be decomposed into successive bifurcations of type $|k|=1$. We believe an analogous statement hold in the discrete setting as well. However, it may require further subdivisions of the simplicial complex $K$. A detailed investigation of this question will be carried out in future work.